PREDISSOCIATION DYNAMICS OF THE A^{2}Σ^{+} STATE OF THE SH RADICAL

\underline{A.J. ORR-EWING} and R.A. ROSE, School of Chemistry, University of Bristol, Bristol BS8 1TS, U.K.; C.-H. YANG, K. VIDMA, and D.H. PARKER, Institute for Molecules and Materials, University of Nijmegen, Toernooiveld 1, 6525 ED Nijmegen, the Netherlands.

Velocity map imaging of the S(^{3}P_{J=0,1,2}) fragments from the predissociation of selected ro-vibrational levels of the A^{2}Σ^{+} state of SH has been used to determine photofragment recoil anisotropies, branching into spin-orbit levels, and the electronic angular momentum polarization of the S atoms. The photofragment recoil anisotropies will be compared with predictions from a model by Houston and coworkers^{a} that accounts for the effects of excited state lifetimes. These vibrational and rotational level dependent lifetimes were previously measured by cavity ring-down spectroscopy and modelled using Fermi Golden rule calculations.^{b} The spin-orbit level branching and populations of m_{J} states, derived from the angular momentum polarization measurements, depend on the correlation of the repulsive states responsible for the A^{2}Σ^{+} state predissociation to the asymptotic H(^{2}S_{1/2}) + S(^{3}P_{J}) limits and non-adiabatic dynamics during the fragmentation.

aH. Kim, K.S. Dooley, S.W. North, G.E. Hall and P.L. Houston, J. Chem. Phys., \underline{125}, 133316 (2006).

bM.D. Wheeler, A.J. Orr-Ewing and M.N.R. Ashfold, J. Chem. Phys., \underline{107}, 7591 (1997).